OpenStudy (anonymous):

clear fractions,solve and check (1/5)(4x - 1) = 7

OpenStudy (anonymous):

try x =9

OpenStudy (anonymous):

\[\frac{1}{5}(4x-1)=7\] \[5\times \frac{1}{5}(4x-1)=7\times 5\] \[4x-1=35\] \[4x=35+1\] \[4x=36\] \[x=\frac{36}{4}\] \[x=9\]